A motorboat can go 8 miles downstream on a river in 20 minutes. It takes 30 minutes for this boat to go back upstream the same 8 miles. Find the speed of the current.
Answer by checkley77(12569)

D=RT
8=R*20
R=8/20
R=2/5 OR .4 MPM. DOWNSTREAM RATE
8=R*30
8=30R
R=8/30
R=4/15 OR .267 MPM. UPSTREAM RATE.
(.4-.267)/2=.133/2=.067 MPM IS THE RATE OF CURRENT.
8=(R+.067)20
8=20R+1.34
20R=8-1.34
20R=6.66
R=6.66/20
R=.333 SPEED OF THE BOAT.
8=(R-.067)30
8=30R-2.01
30R=8+2.01
30R=10.01
R=.333 SPEED OF THE BOAT.